What is the probability that at most two out of 10 customers will prefer the deluxe model


1
Expert's answer
2021-06-14T09:17:07-0400

Let "X=" number of customers who will prefer the deluxe model: "X\sim Bin (n, p)."

A marketing manager makes the statement that the long-run probability that a customer will the deluxe model to the standard model of product is 30%.

Given "n=10, p=0.3, q=1-p=1-0.3=0.7"


"P(X\leq2)=P(X=0)-P(X=1)+P(X=2)"

"=\dbinom{10}{0}(0.3)^0(0.7)^{10-0}+\dbinom{10}{1}(0.3)^1(0.7)^{10-1}"

"+\dbinom{10}{2}(0.3)^2(0.7)^{10-2}"

"=(0.49+2.1+4.05)(0.7)^8\approx0.3828"